In government data, a household consists of all occupants of a dwelling unit, while a family consists of two or more persons who live together and are related by blood or marriage. So all families from households, but some households are not families. Here are the distribution of household size and of family size in the United States:

Number of persons 3 4 5 6 7

Household probability 0.30 0.16 0.03 0.14 0.37

Family probability 0.44 0 0.22 0.09 0.25

Choose an American household at random and let the random variable Y be the number of persons living in the household. What is P(4 < Y < 7)?

0.70

0.31

0.56

0.17
Answer #1

Solution :

From given that ,

P(4 < Y < 7) = P(Y = 5) + P(Y = 6)

= 0.03 + 0.14

= 0.17